A flexible container at an initial volume of 6.13 L contains 4.51 mol of gas. More gas is then added to the container until it reaches a final volume of 16.7 L. Assuming the pressure and temperature of the gas remain constant, calculate the number of moles of gas added to the container.

number of moles of gas: _______ mol
Answer #1

Given:

ni = 4.51 mol

Vi = 6.13 L

Vf = 16.7 L

use:

Vi/ni = Vf/nf

6.13 L / 4.51 mol = 16.7 L / nf

nf = 12.29 mol

Use:

Mol of gas added = nf - ni

= 12.29 mol - 4.51 mol

= 7.78 mol

Answer: 7.78 mol